Gladstorm Enterprises sells a product for $49 per unit. The variable cost is $31 per unit, while fixed costs are $14,256. Determine the following: Round your answers to the nearest whole number.

a. Break-even point in sales units units
b. Break-even point in sales units if the selling price increased to $64 per unit units
Answer #1

a.Contribution margin=Sales-Variable cost

=49-31=$18 per unit

Breakeven=Fixed cost/Contribution margin

=14256/18=792 units

b.Contribution margin=64-31

=$33 per unit

Hence breakeven=14256/33=432 units

Break-even point in sales units 792 units
Break-even point in sales units if the selling price increased to $64 per unit 432 units